# Towards Understanding 3D Vision: the Role of Gaussian Curvature

### v1

@@ 4.1 Middlebury Dataset
We do not report average error for DepthAnythingV2 and DepthPro, as these methods provide only relative depth. Although DepthPro estimates depth in meters and predicts the focal length from an image, our investigation revealed that its depth predictions face real-scale limitations, which is expected for monocular methods. However, to the best of our knowledge, DepthPro provides the most detailed surface estimation among current MDE approaches. While DepthPro tends to minimize GC on the Middlebury dataset (see Figures 6 and 7), both DepthPro and DepthAnythingV2 reconstruct scenes with higher curvatures in our synthetic dataset (see Figure 10). To visualize the results, Figure 12 shows the 3D surface and GC estimated from SOTA MDE approaches, like DepthPro and DepthAnythingV2.
@@ 5 Conclusion
Our evaluation of state-of-the-art methods on the Middlebury benchmark revealed that the GC normalized histograms generated by these methods approximate the high LGC values of the ground truth GC distribution both on the Middlebury dataset and on our 3D synthetic scenes. Then, LGC can be used as a quality measure across multiple 3D reconstruction modalities, including stereo vision, monocular depth estimation, and, by inference, structure from motion. Finally, our work enhances the interpretability and understanding of 3D vision by highlighting Gaussian Curvature as an intrinsic geometric prior for indoor 3D surfaces. Grounded in modern deep learning data, our approach underscores the importance of 3D geometric modeling in capturing critical visual information and can guide the development of next-generation vision systems.
@@ 8.1 Middlebury Dataset
In the main paper, we presented a normalized histogram distribution in Figure 6, which shows the LGC metric for FoundationStereo, DepthPro, RAFT-Stereo, Selective-IGEV, and the ground truth (GT). We selected these techniques for inclusion in the main text because we had already tested their code on our 3D synthetic scenes. Additionally, they represent key categories: FoundationStereo belongs to Group A (new SOTA), RAFT-Stereo and Selective-IGEV to Group B (previous SOTA), and DepthPro is the best-performing monocular depth estimation (MDE) method in terms of depth reconstruction.
@@
Figure 13: Curvature Distribution: each plot presents a normalized histogram of the GC distribution for all 15 training images from the Middlebury dataset. We discarded the highest 20% of $|K|$ values, and plotted the remaining $K$ values within $4,000,4,000]m^{-2}$ in 50 bins uniformly distributed.
@@
Figure 14: Middlebury Ranking: Disparity average error per image (see subsection 8.2 for more information).
Figure 15: Middlebury Curvature: Average Gaussian Curvature per image (see subsection 8.2 for more information).
@@
The results of each aforementioned approach are shown in Figures 16171819, and 20, respectively. For the Group A and Group B approaches, which estimate the disparity of a scene, we computed depth using the equation $Depth=\frac{f\cdot b}{d}$, where $f$ is the focal length in pixels, $b$ is the baseline in meters, and $d$ is the disparity value in pixels. Note that no pixel offset correction is needed ($doffs=0$), as our simulated environment ensures that the left and right images are rectified. The resulting depth for each technique is shown in the first column of its respective figure.
@@
All relevant insights and observations are described in the figure’s captions to support and streamline the reading process.

### v2

@@ 5.1 Middlebury Dataset
We do not report average error for DepthAnythingV2 and DepthPro, as these methods provide only relative depth. Although DepthPro estimates depth in meters and predicts the focal length from an image, our investigation revealed that its depth predictions face real-scale limitations, which is expected for monocular methods. However, to the best of our knowledge, DepthPro provides the most detailed surface estimation among current MDE approaches. While DepthPro tends to minimize GC on the Middlebury dataset (see Figures 6 and 7), both DepthPro and DepthAnythingV2 reconstruct scenes with higher curvatures in our 3D synthetic scenes (see Figure 13). To visualize the results, Figure 15 shows the 3D surface and GC estimated from SOTA MDE approaches, like DepthPro and DepthAnythingV2.
@@
6 Limitations and Future Directions
This paper highlights Gaussian curvature as a powerful measure for analyzing 3D scenes, given its invariance to viewpoint changes and sparsity in man-made environments. However, our analysis is limited to the 15 training images from the Middlebury dataset, as it is the only benchmark that publicly provides the techniques’ results. Although we did not directly evaluate other benchmarks, the strong performance of BLMT-Stereo as the top method on Middlebury [22] and ETH3D [23], and of MonSter++ as the leading method on KITTI 2012 [9] and KITTI 2015 [19], suggests that our findings may generalize to outdoor scenes. As future work, we plan to extend our investigation to additional datasets.
In this paper, we focused on analyzing and understanding the role of Gaussian curvature in 3D reconstruction using SOTA stereo-vision techniques. We did not aim to develop a new algorithm at this stage. As future work, we plan to incorporate low absolute Gaussian curvature as an unsupervised metric for stereo 3D reconstruction, which also includes treating depth discontinuities, occlusions, repetitive patterns, and textureless regions.
@@ 7 Conclusion
Our empirical evaluation of state-of-the-art methods on the Middlebury benchmark revealed that the GC normalized histograms generated by these methods approximate the high LGC values of the ground truth GC distribution both on the Middlebury dataset and on our 3D synthetic scenes. This observation suggests that very likely state-of-the-art methods do incorporate a prior that minimizes GC values in regions where data matching is not sufficient to infer a solution. However, one does not know where in the algorithm such a prior occurs nor, in analogy to modules that perform feature extraction, how to extract a module with such a prior to be used in other applications.
In summary, our work enhances the interpretability and understanding of 3D vision by highlighting Gaussian Curvature as an intrinsic geometric prior for indoor 3D surfaces. Grounded in modern deep learning data, our approach underscores the importance of 3D geometric modeling in capturing critical visual information and can guide the development of next-generation vision systems.
As a possible immediate consequence of this study, LGC could be used as a quality measure in multiple 3D reconstruction modalities, including stereo-vision, monocular depth estimation, and, by inference, structure from motion.
@@ 10.1 Middlebury Dataset
In the main paper, we presented a normalized histogram distribution in Figure 7, which shows the LGC metric for FoundationStereo, DepthPro, RAFT-Stereo, Selective-IGEV, and in Figure 5 for the ground truth (GT). We also present in Table 2 a ranking that includes several benchmarking metrics alongside LGC. We selected these techniques for inclusion in the main text because we had already tested their code on our 3D synthetic scenes. Additionally, they represent key categories: FoundationStereo belongs to Group A (new SOTA), RAFT-Stereo and Selective-IGEV to Group B (previous SOTA), and DepthPro is the best-performing monocular depth estimation (MDE) method in terms of depth reconstruction.
@@
Figure 16: Curvature Distribution: each plot presents a normalized histogram of the GC distribution for all 15 training images from the Middlebury dataset. We discarded the highest 20% of $|K|$ values, and plotted the remaining $K$ values within $5,000,5,000]m^{-2}$ in 50 bins uniformly distributed.
@@
Figure 17: Middlebury Ranking: Average disparity error (AvgError) per image (see subsection 10.1 for more information).
Figure 18: Middlebury Curvature: Average Absolute Gaussian Curvature (Avg|K||K|) per image (see subsection 10.1 for more information).
@@
The results of each aforementioned approach are shown in Figures 19202122, and 23, respectively. For the Group A and Group B approaches, which estimate the disparity of a scene, we computed depth using the equation $Depth=\frac{f\cdot b}{d}$, where $f$ is the focal length in pixels, $b$ is the baseline in meters, and $d$ is the disparity value in pixels. Note that no pixel offset correction is needed ($doffs=0$), as our simulated environment ensures that the left and right images are rectified. The resulting depth for each technique is shown in the first column of its respective figure.
@@
All relevant insights and observations are described in the figure’s captions to support and streamline the reading process. We also omitted some parts of the figures for Blind Review.